### v1

# Future-Proofing Class Incremental Learning

Abstract
Exemplar-Free Class Incremental Learning is a highly challenging setting where replay memory is unavailable. Methods relying on frozen feature extractors have drawn attention recently in this setting due to their impressive performances and lower computational costs. However, those methods are highly dependent on the data used to train the feature extractor and may struggle when an insufficient amount of classes are available during the first incremental step. To overcome this limitation, we propose to use a pre-trained text-to-image diffusion model in order to generate synthetic images of future classes and use them to train the feature extractor. Experiments on the standard benchmarks CIFAR100 and ImageNet-Subset demonstrate that our proposed method can be used to improve state-of-the-art methods for exemplar-free class incremental learning, especially in the most difficult settings where the first incremental step only contains few classes. Moreover, we show that using synthetic samples of future classes achieves higher performance than using real data from different classes, paving the way for better and less costly pre-training methods for incremental learning.
keywords:
Class Incremental Learning , Continual Learning , Image Classification , Image Generation
1 Introduction
Figure 1: Synthetic images of future classes for CIFAR100 generated using different diffusion models. Images are displayed before being resized for better appreciation. First row has been generated by Stable Diffusion 1.4 with a guidance scale of $2.0$, second row using Stable Diffusion 1.4 with a guidance scale of $7.5$, and last row using DALL-E2. From left to right: butterfly, can, house, lion, mountain, orange, otter, pine tree, rose, and streetcar.
While humans naturally learn and accumulate knowledge continuously during their whole lifespan, deep neural networks struggle to do the same. Continuously training deep neural networks on new data will inevitably induce a severe decrease of the performance on the previously learned data. This phenomenon, named catastrophic forgetting [29, 38], has received more and more attention during the past decade, and mitigating it is the core problem of incremental learning and continual learning. Among the different settings of incremental learning, Class Incremental Learning (CIL) is among the most challenging ones. In this setting, the model is incrementally trained on a few number of classes at a time, without having access to the previously learned ones. The objective is to train a unified model that performs well on every encountered class.
The rehearsal memory [6, 39] is one of the most commonly used methods to mitigate catastrophic forgetting and is used by most of the approaches for class incremental learning. It consists of a small buffer containing few representative exemplars of each previously learned class that are later used for replay while learning new classes. However, there remain situations where using a rehearsal memory is not possible, either due to privacy concerns regarding the training data or due to limitations in the storage space. This led to the emergence of a new setting named Exemplar-Free Class Incremental Learning (EFCIL) or Class Incremental Learning Without Memory. Methods for EFCIL often freeze the feature extractor after the first incremental step [13, 35], achieving competitive performances while allowing for a faster and less computationally expensive learning of new classes compared to methods that have to backpropagate through the entire deep neural network. However, this may also lead to poor performance when the number of classes available in the initial step to train the feature extractor is too limited.
In this work, we propose to address this limitation by leveraging pre-trained text-to-image diffusion models to better prepare exemplar-free class incremental learning methods relying on a frozen feature extractor for the future. The diffusion models generate synthetic samples of possible future classes that will be jointly used with actual real data during the first incremental step to train a feature extractor that better generalizes to future classes. Experiments using various challenging settings and datasets for EFCIL highlight how our method can be combined with state-of-the-art approaches and results in significant improvement of the average incremental accuracy. Moreover, it appears that using synthetic images of future classes achieves higher performance than using real images from different yet similar classes.
Our contributions can be summarized into two points:
  • 1.
    We propose a framework relying on pre-trained diffusion models to better prepare exemplar-free class incremental learning methods for the future.
  • 2.
    We experimentally show that using synthetic samples of future classes for training the feature extractor achieves higher performance than using real images from different classes.
2 Related Work
Figure 2: Diagram representing our proposed method Future-Proof Class Incremental Learning (FPCIL) which prepares the feature extractor for the future by jointly training it on an auxiliary dataset in addition to the current dataset during the initial step. The auxiliary dataset contains synthetic samples of future classes generated using a pre-trained text-to-image diffusion model.
2.1 Class Incremental learning (CIL) and exemplar-free class incremental learning (EFCIL)
Most approaches for Class Incremental Learning [3, 28, 60] leverage knowledge distillation [9, 17, 39], a rehearsal memory and a method to mitigate the bias toward the new classes [1, 17, 20, 52] induced by the limited size of the replay memory. Recent works [53, 61] proposed to extend the feature extractor at each incremental step to learn the new classes while preserving past knowledge by freezing previously learned weights.
Exemplar-free class incremental learning is a more challenging setting in which there is no rehearsal memory, making the replay of samples from the previously encountered classes impossible. Similarly to approaches designed for class-incremental learning, numerous methods for exemplar-free class incremental learning also continuously finetune the feature extractor and the classifier [21, 51, 62, 63, 64]. Several authors have highlighted the risk of overfitting on the current task and the importance of having general and transferable feature representation for EFCIL. To address this concern, approaches often rely on self-supervised learning [51, 63] or by creating augmented classes using Mixup [54]. Recently, methods relying on a fixed feature extractor have gained more attention [13, 32]. Those methods either rely on a pre-trained model [45, 49, 50, 57] or by freezing the feature extractor after learning it during the initial step. Petit et al. [35] proposed FeTrIL which achieves state-of-the-art performances by relying on a frozen feature extractor and a linear classifier trained using actual features of new classes and pseudo-features of past classes generated by a geometric translation of new class features. Those methods exhibit numerous advantages including a less computationally expensive training procedure that allows for for a deployment on edge devices. However, they also depend on the initial training strategy used [19, 36] and may achieve poor performance if the quantity of data available for training the feature extractor during the initial step is too limited.
2.2 Training on synthetic images
While few attempts have been made using GAN-based models [4, 18], the use of synthetic images for training deep learning models has recently gained momentum following the advances in diffusion models [15, 46] and their wide availability. Diffusion models generate synthetic samples using an iterative denoising process conditioned on an input, usually a text prompt. Imagen [42] performs the denoising task in the pixel space while Stable Diffusion [40] and DALL-E2 [37] perform it in the latent space in order to decrease the computational cost of the generative process. In one of the pioneer works, Sariyildiz et al. [43] studied the use of the pre-trained text-to-image diffusion model Stable Diffusion to generate a synthetic replacement for the original training dataset. Their experiments highlighted that while the model trained using the real train dataset reaches a higher accuracy on the real test dataset compared to the one trained on synthetic images, they perform on par for transfer learning tasks. Concurrently, Zhang et al. [56] proposed to expand real datasets using synthetic samples generated with various diffusion models and achieved significant improvement on various image classification datasets. Several authors proposed to use synthetic images generated by pre-trained diffusion models for different problems of computer vision such as zero-shot and few-shot learning [14], long-tailed image classification [44], incremental learning [12, 22], out-of-distribution detection [11], or semantic segmentation [31]. Tian et al. [47] showed that the representation learned using contrastive learning on synthetic images generated by Stable Diffusion outperforms state-of-the-art methods trained on real datasets.
2.3 Additional data for Incremental Learning
Several authors [2, 25, 26, 27, 55] have considered the use of additional data for class-incremental learning by leveraging datasets of real images belonging to classes different from the ones that the model has to learn. Most of these approaches assume that the additional dataset is unlabeled, restricting the use of those additional data samples to the distillation loss [26, 27, 55]. Jodelet et al. [22] first proposed to use large pre-trained text-to-image diffusion models in order to generate additional synthetic images for the classes previously encountered by the model. Using a textual prompt based on the class name, this method generates labeled samples that can be used for the distillation loss but also for any supervised losses such as the classification loss. Similarly, pre-trained diffusion models were also used for class-incremental semantic segmentation [7] and object detection [23].
2.4 Preparing for the future
While most of the methods for class-incremental learning only focus on how to prevent the forgetting of past data, some authors have studied how to better prepare the model for future updates. Douillard et al. [10] took inspiration from zero-shot learning and proposed to incorporate into the classifier feature vectors of unseen future classes during training. Those possible future vectors, named ghost features, are produced by a generator using actual features of seen classes and prior information about all the classes. [34] proposed to use a pre-allocated Regular Polytope Classifier which can exploit future unseen classes as negative examples. Zhou et al. [58] proposed to reserve embedding space for future classes by pre-assigning virtual prototypes and by forecasting future classes using manifold Mixup [54]. Bellitto et al.[2] proposed to use an auxiliary dataset of real images containing classes unrelated to the main task in order to prepare the model for the future. Those additional data samples are used to learn more general and stable representation that can be leveraged when learning future classes.
Compared to previously mentioned methods, in this work, we aim to leverage pre-trained text-to-image diffusion models not for generating additional past data but for generating synthetic samples for possible future classes. We study how those synthetic images of future classes can be used during the initial incremental step and to what extent they can help to better prepare exemplar-free incremental learning methods relying on a frozen feature extractor.
Table 1: Definition of the different auxiliary datasets used during the first incremental step for each setting. “Class overlap” is the percentage of classes (actual number of classes in parentheses) in the auxiliary dataset $\mathcal{S}$ that are also in the target dataset, i.e. that will be encountered during the future incremental steps.
Target dataset Auxiliary dataset $\mathcal{S}$ Nature of data Setting Number of classes $\lvert\mathcal{Y}_{\mathcal{S}}\rvert$ Samples per class $n$ Class overlap with target dataset Note
CIFAR100 (100 classes) CIFAR10 Real All 10 5000 0%
FPCIL-Partial (k%) Synthetic B50 Inc 50 500 kkof the classes are randomly selected from the target dataset, the remaining are randomly selected from the wrong predictions of GPT-3.5. For k=100, we use the denomination “FPCIL-Oracle”.
B0 Inc10 90 500 k%
B0 Inc5 95 500 k%
FPCIL-GPT Synthetic B0 Inc10 150 500 35.3% (53) Classes the most often predicted by GPT-3.5 based on the classes encountered in the first incremental step.
B0 Inc5 166 500 30.7% (51)
FPCIL-GPT R1 Synthetic B0 Inc10 90 500 47.8% (43) Subset of “FPCIL-GPT” containing a more restricted set of classes.
B0 Inc5 102 500 31.4% (32)
FPCIL-GPT R2 Synthetic B0 Inc10 53 500 56.6% (30) Subset of “FPCIL-GPT” containing a more restricted set of classes. More restrictive than R1.
B0 Inc5 58 500 34.5% (20)
ImageNet-Subset (100 classes) ImageNet-Compl. Real B50 Inc 50 $\sim$1300 0% Classes randomly sampled from the remaining 900 classes of ILSVRC 2012. No class overlap with ImageNet-Subset.
B0 Inc10 90 $\sim$1300 0%
B0 Inc5 95 $\sim$1300 0%
ImageNet-Compl. X2 Real B50 Inc 100 $\sim$1300 0% Classes randomly sampled from the remaining 900 classes of ILSVRC 2012. No class overlap with ImageNet-Subset.
B0 Inc10 180 $\sim$1300 0%
B0 Inc5 190 $\sim$1300 0%
ImageNet-Compl. Animals Real B0 Inc10 90 $\sim$1300 0% Classes randomly sampled among the “animals” classes from the remaining 900 classes of ILSVRC 2012. No class overlap with ImageNet-Subset.
B0 Inc5 95 $\sim$1300 0%
FPCIL-Oracle Synthetic B50 Inc 50 1300 100% Future classes perfectly predicted.
B0 Inc10 90 1300 100%
B0 Inc5 95 1300 100%
FPCIL-Partial (0%) Synthetic B50 Inc 50 1300 0% Same classes as “ImageNet-Compl.”.
B0 Inc10 90 1300 0%
B0 Inc5 95 1300 0%
FPCIL-Partial Animals (0%) Synthetic B0 Inc10 90 1300 0% Same classes as “ImageNet-Compl. Animals”.
B0 Inc5 95 1300 0%
3 Proposed method
The objective of Class-Incremental Learning (CIL) is to learn a unified model $\theta$ on a gradually increasing set of classes. $\theta$ can be decomposed into $\{\Phi,\mathcal{W}\}$ with $\Phi$ the feature extractor and $\mathcal{W}$ the classifier. The training procedure is divided into $T$ incremental steps, each step consisting of a dataset $\mathcal{D}_{t}$ containing samples from the set of new classes $\mathcal{Y}_{t}$ such that $\forall i,j\in\{1,2,...,T\},\mathcal{Y}_{i}\cap\mathcal{Y}_{j}=\emptyset$ for $i\neq j$. Contrary to standard Class-incremental Learning, in the Exemplar-Free Class Incremental Learning (EFCIL) setting, the model does not have access to any replay memory: during each incremental step $t$, the model $\theta_{t}$ is trained solely on the current dataset $\mathcal{D}_{t}$. After each incremental step, the model is evaluated on the test dataset of all the classes observed so far without having any step descriptor.
3.1 Future-Proof Class Incremental Learning
In order to create a future-proof model, as illustrated in Figure 2, we introduce Future-Proof Class Incremental Learning (FPCIL), a straightforward approach which can be combined with any method for class incremental learning relying on a frozen feature extractor. During the first incremental step, when training the feature extractor, in addition to the initial step dataset $\mathcal{D}_{1}$, an additional dataset $\mathcal{S}$ containing $\lvert\mathcal{Y}_{\mathcal{S}}\rvert$ different classes is used. While any labeled datasets could be used for $\mathcal{S}$, we propose a novel approach which consists in using images from classes that will be encountered by the model during the future incremental steps. In detail, the objective of the initial incremental step is extended from solely learning the initial classes $\mathcal{Y}_{1}$ into learning $\mathcal{Y}_{1}\cup\mathcal{Y}_{\mathcal{S}}$ where $\mathcal{Y}_{\mathcal{S}}=\bigcup_{i=2}^{T}\mathcal{Y}_{i}$ are the future classes. The feature extractor is jointly trained with a classifier on $\mathcal{D}_{1}\cup\mathcal{S}$. At the end of the initial step, the classifier is restricted to the base classes $\mathcal{Y}_{1}$ by dropping the weights corresponding to the future classes from the auxiliary dataset $\mathcal{S}$, and the feature extractor is frozen. Then, the training procedure continues according to the baseline method for the remaining incremental steps.
In the initial step, our proposed method FPCIL trains on both the initial training dataset $\mathcal{D}_{1}$ and an additional dataset $\mathcal{S}$ which contains training data for predicted future classes. As only the data for the current step are available in the Class Incremental Learning setting, we propose to use synthetic data instead of the real ones. We use a pre-trained text-to-image diffusion model to generate synthetic images from the future classes during the initial step. This means that only the labels of the future classes are necessary. We argue that this requirement is practical and is compatible with most real-world applications of class incremental learning. For example, when initially creating a deep learning model, it is common to plan to update the weights later as new data are collected. The nature of the future data is generally known, in particular future classes that may appear, as the data collection process has to be planned and designed in advance due to its high cost in time and resources. More generally, when deploying a learning agent, realistic assumptions on the classes that will be encountered or not can be made as the agent evolves together with the environment; that can be used to prepare the learning agent accordingly.
Based on our experiments detailed in the ablation study, the classes in the auxiliary dataset $\mathcal{S}$ do not need to be perfectly predicted to have a positive impact on the performance. For clarity, we refer to FPCIL as “FPCIL-Oracle” when the list of future classes has been perfectly predicted and “FPCIL-Partial (k%)” otherwise, with $k$ the percentage of future classes correctly predicted.
Predictions of the future classes can be done through expert knowledge or through automated methods. Here, we demonstrate an example of automatically producing predictions for future classes based on the recently popular large language models. Our proposed method relies on the GPT model [5] “GPT-3.5 Turbo Instruct” used through the OpenAI API. The large language model is queried for a completion task using the prompt “The dataset contains the following 100 classes: $c_{1}$, […], $c_{k}$,” where $c_{1}$ to $c_{k}$ are the names of the $k$ classes from the initial step, with $k$ equals to $\lvert\mathcal{Y}_{1}\rvert$. The number “100” is used here to encourage the model to generate a longer list of possible future classes as initial experiments without specifying a number would result in only a few number of guesses. In practice, the output of GPT has to be parsed to extract the predicted future classes. The completion function of GPT often continues the prompt as if it was part of a website or a scientific paper by appending imaginary details such as the number of images per class in the dataset, nonexistent download url for the dataset, or architecture of the model. Due to the random nature of the output from GPT, the prompt is queried 10 times and the predicted classes are ordered according to the number of times they have been predicted. The auxiliary dataset $\mathcal{S}$ is then restricted to the classes which were the most often predicted. We evaluated three versions with an increasing level of restriction: “FPCIL-GPT”, “FPCIL-GPT R1”, and “FPCIL-GPT R2”.
Table 2: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
Methods CIFAR100 - LFH CIFAR100 - LFS
B50 Inc10 B50 Inc5 B50 Inc2 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last AVG Last AVG Last
FeTrIL [35] 68.05 58.69 66.81 58.22 65.46 56.94 50.50 33.86 40.80 24.65
FeTrIL w/ aux. CIFAR10 70.60 60.82 69.34 60.03 68.23 59.26 58.56 41.72 52.42 35.03
FeTrIL w/ FPCIL-Oracle 73.44 65.71 72.81 65.20 72.40 64.76 68.09 55.43 62.92 49.67
FeTrIL w/ FPCIL-Partial (0%) 70.25 61.41 69.27 60.56 68.13 59.67 62.63 47.71 57.46 42.71
FeTrIL w/ FPCIL-GPT - - - - - - 67.18 53.89 62.15 48.67
FeTrIL w/ FPCIL-GPT R1 - - - - - - 66.79 53.02 60.65 46.08
FeTrIL w/ FPCIL-GPT R2 - - - - - - 65.25 51.01 59.89 44.33
Table 3: Performances on ImageNet-Subset with the Learning from Half and Learning From Scratch procedures. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
Methods ImageNet-Subset - LFH ImageNet-Subset - LFS
B50 Inc10 B50 Inc5 B50 Inc2 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last AVG Last AVG Last
FeTrIL [35] 72.97 64.49 72.01 63.54 71.02 62.19 51.64 34.53 39.64 22.42
FeTrIL w/ aux. ImageNet-Compl. 75.19 67.81 74.08 66.96 73.25 65.71 70.62 59.53 68.14 57.45
FeTrIL w/ aux. ImageNet-Compl. X2 76.33 69.52 75.30 68.53 74.69 67.89 73.96 64.68 73.01 62.98
FeTrIL w/ FPCIL-Oracle 77.21 71.80 76.81 71.51 76.84 71.17 74.04 65.65 72.64 64.29
FeTrIL w/ FPCIL-Partial (0%) 74.90 67.27 73.93 66.37 73.18 65.25 68.89 57.21 65.94 55.09
3.2 Generating future data
The synthetic images dataset of future classes is generated following the procedure initially defined by Sariyildiz et al.[43].
The auxiliary dataset $\mathcal{S}$ is generated using a pre-trained diffusion model $G$ conditioned solely on a textual prompt. As proposed in [43], the default prompt used is $c,d_{c}$ with $c$ the name of the class and $d_{c}$ a simple description of the class. Adding the description of the class into the prompt is intended to ensure that the generated images belong to the targeted class and not to any homonyms: for example “crane” in the ImageNet dataset, which may refer to either the bird or the machine used to move heavy objects. The definition $d_{c}$ also tends to restrain the domain of the generated images even though the domain could also be enforced by adding it to the prompt. To automatically create the prompts, we rely on WordNet [30] and use the lemmas of the synset as $c$ and the definition of the synset as $d_{c}$. For each class, $n$ synthetic images are generated with $n$ equals to the number of images per class in the target dataset. Compared to traditional pre-training and concurrent method [2], our proposed method does not require the costly and fastidious process of collecting, curating, and annotating a dataset of real images.
The guidance scale is an important hyper-parameter of diffusion models relying on classifier-free guidance [16]. To some extent, it can be used to control the trade-off between the diversity and the quality of the generated synthetic images. For Stable Diffusion, the default recommended value for the guidance scale is often $7.0$ but following [22, 43] we use the lower value of $2.0$ as a higher diversity is beneficial for pre-training.
4 Experiments
4.1 Experimental setup
4.1.1 Datasets
The experiments are conducted on two large scale datasets: CIFAR100 [24] and ImageNet-Subset. CIFAR100 is a dataset containing 60,000 32x32 RGB images evenly divided among 100 classes with 500 training images and 100 testing images per class. ImageNet-Subset is a subset of the high-resolution images dataset ImageNet (ILSVRC 2012) [41]. This subset contains 100 classes with about 1,300 training images and 50 testing images per class. Following Rebuffi et al. [39], the classes are ordered using numpy with the random seed 1993.
Two training procedures are used: Training from Scratch (TFS) and Training From Half (TFH). For the Training from Scratch procedure, the classes are evenly divided into $T$ (10, or 20) incremental steps, each incremental step containing the same number of classes. For the Training From Half procedure, following [17], the first step contains half of the classes and the remaining classes are evenly divided into the following $T-1$ incremental steps. Following Zhou et al. [60], the two training procedures are referred to using the common naming system “B-$i$ Inc-$j$ with $i$ the number of classes in the first step and $j$ the number of classes each of the following incremental step; $i$ equals to zero is used to refer to the Learning From Scratch procedure.
4.1.2 Model training
Following common practice for exemplar-free class incremental learning [35, 63, 64], ResNet-18 is used for both CIFAR100 and ImageNet-Subset. The feature extractor is trained solely during the first incremental step and is frozen for the rest of the training procedure. The backbone is trained with SDG using the standard Softmax Cross-Entropy loss and AutoAugment [8] for 160 epochs with a batch size of 128 and a weight decay of $5e^{-4}$. The initial learning rate is set to 0.1 and decrease to zero using a cosine annealing schedule. For the classifier, we use a fully-connected linear layer trained following FeTrIL [35]. The classifier is trained using SGD for 50 epochs at each incremental step, with an initial learning rate of 0.1 and cosine annealing schedule. In the ablation study, we also report results using a Nearest Mean Classifier (NMC). Experiments are conducted using Pytorch [33] and PyCIL [59].
4.1.3 Synthetic generation
Two diffusion models are used for generating the synthetic samples during training: Stable Diffusion [40] and DALL-E2 [37]. If not specified otherwise, the default diffusion model is Stable Diffusion v1.4  [^1] https://huggingface.co/CompVis/stable-diffusion-v1-4 used with a guidance scale of $2.0$ for 50 steps. Following [43], $c,\;d_{c}$ is used for the prompt. The same datasets of synthetic images are used for all the experiments. Those datasets have been generated prior to incremental learning experiments, at size 512x512 using a fixed seed and contains 500 images per class for CIFAR100 and 1300 images per class for ImageNet-Subset. For CIFAR100, the synthetic images are resized to 32x32 before being fed to the model. For DALL-E2, because the model is used through the official API of OpenAI, it is not possible to control the different hyper-parameters of the model; the default, unknown, values are used for the guidance scale, the random seed and the number of denoising step. A slightly different prompt is used for DALL-E2, “a photo of a $c,\;d_{c}$, as the default one would generate unsatisfactory images for some classes. In order to comply with the moderation policy of the API, few definitions $d_{c}$ have been changed.
4.1.4 Baselines
As a comparison baseline, a dataset of real images containing general and/or similar classes to the ones in the target dataset is used as auxiliary dataset during the first incremental step. This baseline is adapted from the method proposed by Bellitto et al.[2] and applied to exemplar-free class incremental learning with a frozen feature extractor. For CIFAR100, we use CIFAR10 for the auxiliary dataset as both datasets were created using the same annotation and curation methodology from the TinyImages Dataset [48]. However, while CIFAR10 contains more samples than the datasets generated using our proposed method FPCIL, it is important to note that CIFAR10 contains only 10 classes. For ImageNet-Subset, the auxiliary dataset “ImageNet-Compl.” was created by randomly selecting classes from the complement of ImageNet-Subset in ImageNet. “ImageNet-Compl.” contains the same number of classes as the datasets generated by our proposed method FPCIL and approximatively the same number of samples. This is a strong baseline as both subsets are sampled from the same dataset and, while they do not contain the exact same classes, classes in both datasets are from the same categories and some are highly similar: for example both datasets contain different breeds of dogs or species of insects. We additionally use “ImageNet-Compl. X2” which contains two times more classes than ImageNet-Compl. We also report the performance of “FPCIL-Partial (k%)” which generates a dataset of the same size as FPCIL-Oracle but with only $k$of the classes correctly predicted. For CIFAR100, this auxiliary dataset is built by randomly sampling $k$of the classes from FPCIL-Oracle and the remaining ones from the wrong prediction of GPT-3.5. For ImageNet-Subset, “FPCIL-Partial (0%)” contains synthetic images belonging to the same classes as the ones in ImageNet-Compl. The definition of the all different auxiliary datasets used are summarized in Table 1.
4.1.5 Evaluation
Models are evaluated and compared using both the “Final Accuracy” and the “Average Incremental Accuracy” [39]. The Average Incremental Accuracy is defined as the average of the Top-1 accuracy of the model on the test dataset of all the classes learned so far, at the end of each incremental step including the first one.
4.2 Results
Average incremental accuracy and final accuracy of our proposed method FPCIL are reported for CIFAR100 and ImageNet-Subset in Table 2 and Table 3 for both the Learning From Scratch and Learning From Half procedures.
Table 4: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures for FeTrIL and Nearest Mean Classifier (NMC). Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
Methods CIFAR100
B50 Inc5 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last
FeTrIL [35] 66.81 58.22 50.50 33.86 40.80 24.65
FeTrIL w/ FPCIL-Oracle 72.81 65.20 68.09 55.43 62.92 49.67
   Improvement in p.p. +6.00 +6.98 +17.59 +21.57 +22.12 +25.02
NMC 67.35 57.08 40.28 23.49 29.55 14.34
NMC w/ FPCIL-Oracle 73.57 65.31 62.13 47.81 57.13 42.53
   Improvement in p.p. +6.22 +8.23 +21.85 +24.32 +27.58 +28.19
Table 5: Performances on CIFAR100 with the Learning From Scratch procedure depending on the ratio of correctly predicted future classes in the complementary dataset used during the initial step. Synthetics samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
Methods CIFAR100 - B0 Inc10
AVG Last
FeTrIL [35] 50.50 33.86
FeTrIL w/ FPCIL-Oracle (100%) 68.09 55.43
FeTrIL w/ FPCIL-Partial (66%) 66.06 52.21
FeTrIL w/ FPCIL-Partial (50%) 64.79 50.76
FeTrIL w/ FPCIL-Partial (33%) 64.32 49.90
FeTrIL w/ FPCIL-Partial (0%) 62.63 47.71
Table 6: Performances on ImageNet-Subset with the Learning From Scratch procedure depending on the complementary dataset used during initial step. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
Methods ImageNet-Subset - LFS
B0 Inc10 B0 Inc5
AVG Last AVG Last
FeTrIL [35] 51.64 34.53 39.64 22.42
FeTrIL w/ FPCIL-Oracle 74.04 65.65 72.64 64.29
FeTrIL w/ aux. ImageNet-Compl. Animals 68.74 57.87 64.09 54.29
FeTrIL w/ aux. ImageNet-Compl. 70.62 59.53 68.14 57.45
FeTrIL w/ FPCIL-Partial Animals (0%) 66.08 54.97 61.63 51.11
FeTrIL w/ FPCIL-Partial (0%) 68.89 57.21 65.94 55.09
Our proposed approach FPCIL considerably improves the average incremental accuracy of the state-of-the-art method FeTrIL [35], from 5.39 percentage points (p.p.) up to 22.12p.p. on CIFAR10 and from 4.24p.p. up to 33.00p.p. on ImageNet-Subset. This improvement is notably higher for the more difficult Learning From Scratch procedure which is particularly challenging for FeTrIL due to the limited quantity of data available during the first incremental step. Compared to using an auxiliary dataset [2] of real images from classes different from the ones in the target dataset, our method also achieves significantly higher accuracy. On ImageNet-Subset, our proposed method FPCIL even outperforms a two times larger auxiliary dataset containing real images. Moreover, even in the case of only wrong guesses for the future classes, the auxiliary dataset of synthetic images performs on par with concurrent approaches relying on real datasets: for the Learning From Half procedure, the difference of average incremental accuracy between using ImageNet-Compl or its synthetic counterpart is at most 0.29p.p.. Furthermore synthetic datasets can be rapidly generated whereas auxiliary datasets of real data require a long and costly collection, curation, and annotation process. All those points highlight the core of our proposed method: while simply using additional real data for pre-training the feature extractor is beneficial, using data from future classes instead of random ones is better, even if the data used are synthetic.
On CIFAR100, Our more practical approach FPCIL-GPT also reaches a higher accuracy than the baseline method FeTrIL yet slightly lower than the FPCIL-Oracle. On CIFAR100 in the Learning From Scratch procedure with 10 steps of 10 classes (B0 Inc10), the less restricted FPCIL-GPT predicted a total of 150 classes with 53 (35,3%) that actually are among the 90 future classes while the most restricted FPCIL-GPT R2 predicted only 53 classes with 30 (56.6%) among the 90 future classes. All the details regarding future class predictions using GPT are reported in Table 1. The more restricted version FPCIL-GPT R1 and R2, tend to have a higher ratio of correctly predicted future classes but also a lower total number of predicted classes, resulting in a lower accuracy than the less restricted FPCIL-GPT. This emphasizes that both the correctness of the prediction of future classes and the actual quantity of data used for pre-training are important. In the ablation study, we further analyze the impact of the correctness of the prediction of future classes.
4.3 Ablation study
Table 7: Performances on CIFAR100 with the Learning from Half and Learning From Scratch procedures depending on the Diffusion Model used for generating the synthetic data: Stable Diffusion with a guidance scale of $2.0$, Stable Diffusion with a guidance scale of $7.5$ or DALL-E2. “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
Methods CIFAR100
B50 Inc5 B0 Inc10 B0 Inc5
AVG Last AVG Last AVG Last
FeTrIL [35] 66.81 58.22 50.50 33.86 40.80 24.65
FeTrIL w/ FPCIL-Oracle SD $g=2.0$ 72.81 65.20 68.09 55.43 62.92 49.67
FeTrIL w/ FPCIL-Oracle SD $g=7.5$ 72.22 64.40 66.48 53.36 61.63 48.50
FeTrIL w/ FPCIL-Oracle DALL-E2 71.08 63.02 63.94 50.39 59.76 46.34
Table 8: Performances on CIFAR100  8a) B50 Inc5 and  8b) B0 Inc10 depending on the backbone used. Synthetic samples generated using Stable Diffusion with a guidance scale of $2.0$ “AVG” is Average Incremental Accuracy (Top-1) and “Last” is the final overall accuracy of the model after the last incremental step. Results averaged over 3 random runs.
Methods CIFAR100 - B50 Inc5
ResNet-32 ResNet-18
AVG Last AVG Last
FeTrIL [35] 64.27 55.19 66.81 58.22
FeTrIL w/ aux. CIFAR10 64.48 55.85 69.34 60.03
FeTrIL w/ FPCIL-Oracle 66.24 58.88 72.81 65.20
FeTrIL w/ FPCIL-Partial (0%) 63.80 55.67 69.27 60.56
(a) CIFAR100 - B50 Inc5
Methods CIFAR100 - B0 Inc10
ResNet-32 ResNet-18
AVG Last AVG Last
FeTrIL [35] 46.32 29.80 50.50 33.86
FeTrIL w/ aux. CIFAR10 56.70 39.42 58.56 41.72
FeTrIL w/ FPCIL-Oracle 62.73 48.96 68.09 55.43
FeTrIL w/ FPCIL-Partial (0%) 59.68 45.30 62.63 47.71
(b) CIFAR100 - B0 Inc10
4.3.1 Accurately predicting the future
In Table 5, we compare the performance of our proposed methods combined with FeTrIL on CIFAR100 B0 Inc10 depending on the ratio of correctly predicted future classes in the auxiliary dataset used during the initial step. Experimentally, for a fixed dataset size, the average incremental accuracy decreases with the correctness of the prediction of the future classes. This confirms the advantage of predicting the future classes compared to using random classes which may or may not be related to the ones that will be encountered by the continual learner.
Additionally, in Table 6 we report the accuracy for an alternative version of ImageNet-Compl and FPCIL-Partial named respectively “ImageNet-Compl. Animals” and “FPCIL-Partial Animals”. ImageNet-Compl Animals is built by sampling classes from the complement to ImageNet-Subset in ImageNet while restricting the classes to animals; FPCIL-Partial Animals contains the same classes that ImageNet-Compl Animals but the images are generated using a diffusion model. These two more specialized datasets achieve notably lower accuracy than their more general version. This further emphasizes the importance of predicting the future classes and not limiting the scope of the auxiliary dataset.
4.3.2 Architecture of feature extractor
Table 8 compares the performance of our method for two different feature extractors: the 11.2 million parameters ResNet-18 and the 0.46 million parameters ResNet-32. Due to the addition of the auxiliary dataset to the often limited dataset of the initial step, our method takes advantage of a larger feature extractor and results in a large improvement with the baseline compared to backbones with less parameters. Nonetheless, FPCIL remains competitive with smaller deep neural networks.
4.3.3 Other classifiers
Additionally to FeTrIL [35], we report in Table 4 the performance of our method combined with a Nearest Mean Classifier (NMC). It is a simple yet effective method [19, 32] for class incremental learning with a fixed feature extractor that does not require memory.
Similarly to FeTrIL, the performance of NMC is highly dependent on the quality of the features produced by the backbone. It struggles with the Learning From Scratch procedure where the number of classes and data samples used to train the feature extractor is limited. Our method improves the performance of the NMC in every considered setting and achieves the most significant improvements with the Learning From Scratch procedure by reaching up to three times the final accuracy of the baseline method on CIFAR100 B0 Inc5.
4.3.4 Quality and diversity
In Table 7, the performance of FPCIL Oracle is reported on CIFAR100 depending on the diffusion model used to generate the synthetic images of future classes. By default in our study, we used Stable Diffusion with a guidance scale of $2.0$ [43] to favor the diversity to the detriment of the quality. We verified this assumption by comparing it against Stable Diffusion with a higher guidance scale and DALL-E2 which both tend to favor more the quality than the diversity. Experiments confirm that when pre-training the feature extractor on future classes, the diversity is more important than the quality of the generated samples: Stable Diffusion with a small guidance scale achieves the highest results, ahead of DALL-E2 and Stable Diffusion with the default guidance scale.
5 Conclusion
In this work, we introduced a new method for Exemplar-Free Class Incremental Learning which leverages large pre-trained diffusion model to generate images from future classes. Experimental results highlight how our method can significantly improve the accuracy of state-of-the-art approaches while only modifying the initial step. We found that our method requires less data than traditional methods relying on real curated datasets. While we limit our present study to feature extractors trained from scratch during the first incremental step, in future work, we will further investigate how synthetic images of future classes can be used to adapt a general pre-trained foundation model. Moreover, we will also explore how synthetic data of future classes can be leveraged during the whole training procedure, not only the first incremental step.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgment
This work is partly supported by JSPS Grant-in-Aid for Scientific Research (grant number 23H03451, 21K12042) and the New Energy and Industrial Technology Development Organization (Grant Number JPNP20017).